I need to compare interest rates to find the best one, but I am unsure of the formula to use.
For example (1)25 year loan for $390000 with interest 7.4004% compounded monthly
(2) 30 year loan for $390000 with interest 6.76% compounded weekly
(3)20 year loan $390000 with interest 7.2% compounded quartely
Any help would be much appreciated

Answer by stanbon(75887)

I need to compare interest rates to find the best one, but I am unsure of the formula to use.
----
monthly payment = (rP)/[1-(1+r)^(-N)
r is period interest rate
P is amt. borrowed
N = total # of periodic payments
--------------------------------------
For example
(1)25 year loan for $390000 with interest 7.4004% compounded monthly
---
mp = [(0.074004/12)*390000)/[1-(1+(0.07004/12))^(-25*12) = $2913.49
-----------------------------------------------
(2) 30 year loan for $390000 with interest 6.76% compounded weekly
weekly payment = [(0.0676/52)*390000]/[1-(1+(0.0676/52))^(-52*30) = $583.95
So monthly payment would be (52/12)(583.95) = $2530.44
-----------------------------------------------------
(3)20 year loan $390000 with interest 7.2% compounded quartely
Quarterly payment = [(0.072/4)*(390000)]/[1-(1+(0.072/4))^(-4*20) = $9236.62
Monthly payment = (4/12)(9236.62) = $3078.87
================================================================
Cheers,
Stan H.
================